OpenStudy (tainted):

Solve x2 = 14x – 49 (showwork for medal)

OpenStudy (anonymous):

it is easy man x2 = 14x – 49 = x2 - 14x + 49 =0 = (x-7)^2=0 = x=7

OpenStudy (babystar16):

is that suppose to be x to the second power or 2x